OpenStudy (anonymous):

PLEASE HELP WILL GIVE MEDAL! Question 3. 3. Solve the following equation by transforming it into a perfect square trinomial. x2 - 14x = 15 (Points : 1) {57, 71} {-22, 36} {-15, 29} {-1, 15} Question 4. 4. Solve the following equation by transforming it into a perfect square trinomial. 3x2 + 18x = 21 (Points : 1) {-7, 1} {-19, 13} {-25, 7} {-13,- 5}

OpenStudy (solomonzelman):

14= 7×2 and 7²=49

OpenStudy (nincompoop):

do you want to upload all of your homework for the entire semester?

OpenStudy (solomonzelman):

LOL !

OpenStudy (anonymous):

Okay i don't need sass thanks

OpenStudy (solomonzelman):

I would be willing to do it, how much are you paying ? 25cent per problem. All steps and explanation guaranteed.

OpenStudy (anonymous):

25 it is lol

OpenStudy (solomonzelman):

hehe

OpenStudy (anonymous):

\[3 x^2+18x=21 ~or ~3x^2+18 x-21=0\] \[make ~the~coefficient~of~x^2~as~1~by~dividing~by~3\] x^2+6x-7=0 now add both sides\[\left( \frac{ coefficient~of~x }{ 2 } \right)^2~i.e.,\left( \frac{ 6 }{ 2 } \right)^2=9\] \[x^2+6 x+9=7+9\] \[\left( x+3 \right)^2=4^2,x+3=\pm~4\] solve for x.

OpenStudy (anonymous):

Thank you but I got it figured out. :) thanks a lot.

OpenStudy (anonymous):

correction \[3 x^2+18x=21~or~x^2+6x=7\]

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

:)